# Debias your Large Multi-Modal Model at Test-Time with Non-Contrastive Visual Attribute Steering

### v1

@@ Abstract
Large Multi-Modal Models (LMMs) have demonstrated impressive capabilities as general-purpose chatbots that can engage in conversations about a provided input, such as an image. However, their responses are influenced by societal biases present in their training datasets, leading to undesirable differences in how the model responds when presented with images depicting people of different demographics. In this work, we propose a novel debiasing framework for LMMs that directly removes biased representations during text generation to decrease outputs related to protected attributes, or even representing them internally. Our proposed method is training-free; given a single image and a list of target attributes, we can ablate the corresponding representations with just one step of gradient descent on the image itself. Our experiments show that not only can we can minimize the propensity of LMMs to generate text related to protected attributes, but we can improve sentiment and even simply use synthetic data to inform the ablation while retaining language modeling capabilities on real data such as COCO or FACET. Furthermore, we find the resulting generations from a debiased LMM exhibit similar accuracy as a baseline biased model, showing that debiasing effects can be achieved without sacrificing model performance.
@@ 1 Introduction
Deep neural networks are well known to exhibit a myriad of societal biases learned from their training datasets [8, 45]. While this effect is true to some extent for most models, it is especially prevalent in models trained on web-scale data. Numerous prior works have observed and identified such biases present in modern Large Language Models (LLMs) [5, 9], while other recent work shows that such biases are even more prevalent in Large Multimodal Models (LMMs)  [^1] LMMs refer specifcally to generative multimodal models. We refer to the general class of vision-language models as VLMs [7] such as LLaVA-1.5 [24], or Llama-3.2 [12] that integrate a vision backbone with a pretrained LLM. Given that LLMs are often pretrained on relatively uncurated web-scale data [32], any further integration with additional models will inherit the particular biases of the LLM. Without additional safety tuning, these pre-existing biases may be amplified further when an LLM is augmented with pretrained visual capabilities, which also come with a distinct set of implicit societal biases present in the visual pretraining data. Evaluating and mitigating potentially harmful behaviors induced by these societal biases is becoming increasingly important in order to safely deploy multimodal generative AI systems that utilize LMMs. In this work we focus on reducing the propensity of an LMM to attend to protected attributes unnecessarily. While it is often benign and even useful to make use of protected attributes in generation, LMMs have been found to suffer from group-dependent sentiment [27], well as conditional inaccuracies given certain protected attributes [18, 44].
Figure 1: An overview of our model steering approach for LMMs. By solving for an appropriate steering vector we can augment the generation patterns of the LMM by ablating the internal representation corresponding to a given attribute at test time. The steering vector is computed with a single step of gradient optimization to maximize the likelihood of protected tokens. The attribute that we ablate is then computed by taking a difference between original activations and optimized activations.
@@
Recently, a variety of methods have been proposed for debiasing LLMs and Discriminative Vision Language Models individually [22, 34]. However, relatively little prior work has focused specifically on debiasing LMMs. One naive approach could be to simply define a list of words or symbols that are blocked at inference time. While a block-list could have merit in some applications, a more nuanced approach is required for attributes with many or ambiguous meanings, such as perceived race (e.g., black, white) [^2] Wherever we use words such as race or gender throughout this work, we mean perceived race and perceived gender. The labels we used for such social attributes are inherited from prior work and are not a reflection of our own value judgments. See Section 6 for additional discussion.. Many other existing debiasing approaches for LLMs and VLMs focus on fine-tuning models with additional curated data to reduce bias. However, training is a blunt tool that often results in undesirable outcomes such as a degradation in task-specific performance. Fine-tuning is also labor and computationally intensive, requiring the collection of an additional (likely large-scale) dataset that can appropriately debias the model. Despite prior efforts [17], there remains no cannonical recipe for constructing a dataset large enough to meet data requirements for training LMMs, while also fine-grained enough to ablate specific target attributes like perceived race. In contrast, our work introduces a training-free approach to debiasing LMMs that can be applied to any attribute at inference time without degrading model performance.
@@
Inspired by mechanistic interpretability and prior work in steering language models, we propose a novel method for identifying and ablating biased representations present in LMMs. Recently, prior work has shown that it is possible to ablate representations with respect to target attributes by using a small number of contrastive prompts [26] in order to identify directions in representation space that correspond to an increased propensity to generate related outputs. Our method differs from prior LLM steering work in two fundamental ways. First, our method targets multimodal models rather than just language models. Second, our approach requires no additional data other than an input image and a target attribute to remove. We find that we can leverage the additional information available to a LMM, the input image, in order to solve for the steering direction on the fly. Specifically, during inference we perform a single step of gradient descent on the visual representation in the direction of the target attributes. This one-step optimization is similar to the fast sign gradient method (FGSM) [15] applied on LMM representations. We then can perform model steering as normal, by extracting downstream activations that follow from both the perturbed and the unchanged visual representations. The normalized difference between them is used as the steering direction.
@@
  • We propose an efficient training-free algorithm to find effective steering directions for LMMs that performs a single optimization step on the input visual representation.
  • We show that our proposed approach can effectively debias two popular LMM model families LLaVA and Llama3. We reduce their propensity to generate text related to protected attributes on the SocialCounterfactuals, COCO, and FACET datasets. We find that our results hold across perceived race, physical appearance, age, and gender attributes.
  • We analyze the steered and unsteered generations across all datasets and find that steering does not degrade accuracy or language modeling capabilities, finding instead, steering produces equitable group-level sentiment.
@@ 2 Related Work
Model Steering via Interventions. Due to the expense and difficulty of fine-tuning LLMs for arbitrary domains, its natural to seek an efficient and fine-grained approach for controlling LLM outputs. Prior work has attempted to steer generations by intervening on inputs, outputs, or intermedate representations. Modifying prompts to induce or ablate a feature [41, 40] is one such approach, but is commonly an iterative and inefficient method. Intervening on outputs has been explored via token-level block lists [14], domain classifiers [21] to trigger regeneration, or decoding strategies that decrease probabilities of generating certain tokens [31]. Intervening on intermediate representations is most relevant to our approach. This often entails the discovery of interpretable features that not only explain model behavior [47, 19], but can also be used to intervene and steer the model towards output generations with certain characteristics or content.
Templeton et al. [36] showed that this can be achieved by applying a sparse autoencoder to decompose the activations of an LLM into separable features. There, the authors demonstrated the existence of monosemantic features that can trigger relevant downstream behavior or content when manually introduced during inference. Turner et al. [38] showed LLMs can be steered with a simple contrastive prompt using vector addition, and Arditi et al. [3] demonstrated that the refusal behavior in LLMs can be suppressed through a single vector which can be learned via ablation on representative data and applying a difference-in-means [4] approach. Various methods of steering have also been applied to toxiciy [25] and other behaviors such as hallucination [28].
@@
While several approaches have been proposed for mitigating social biases in VLMs [39, 6, 43, 33, 11, 35, 18], prior research on addressing such biases in LMMs is lacking. Sathe et al. [30] and Fraser and Kiritchenko [13] utilized synthetically generated images to analyze the presence of bias in LMMs, but do not address bias mitigation strategies. Howard et al. [17] also leveraged synthetic images from the SocialCounterfactuals dataset [18] to measure bias in LMMs but at a much larger scale, finding that LMMs possess more bias than the corresponding LLM from which they were trained. They also investigated the usefulness of prompting strategies to reduce bias at inference time, but found that it produced inconsistent debiasing effects across different models and generation settings. While representaiton-based steering for reducing societal biases has been demonstrated in LLMs such as Claude 3 [36], our work is the first to demonstrate successful inference-time representation steering for reducing bias in LMMs.
Figure 2: A detailed description depicting our single step gradient-based steering direction computation. (Step 1) Compute the intermediate activations of the LLM when asking the model a direct question about the target attribute to remove. (Step 2) Run a single step of gradient optimization with a logit maximizing loss function in order to compute a modified version of the image tokens $\mathbf{u}$ which increases the probability the model will reference the protected attribute. (Step 3) Recompute the activations of the LLM with the new modified image tokens $\mathbf{u^{\prime}}$ and take a difference between the two intermediate LLM activations. (Step 4) Run the intended query for the LMM, but ablate the attribute $a$ from all intermediate layers of the LLM.
@@ 3 Methodology
Our approach to debiasing LMMs involves identifying and ablating the bias attribute in the model’s internal representations. We achieve this by contrasting the model’s original activations from alterated activations: either from a single step gradient optimized version or from a collection of bias invoking questions. An overview of the single step gradient-based method is shown in figure 2.
@@ 3.1 Steering Direction Estimation
Dataset-based Estimation
Let $\mathcal{M}$ denote an arbitrary LMM, and $\mathbf{h}^{l}\in\mathbb{R}^{d}$ represent the activations at layer $l$, where $d$ is the dimensionality of the hidden state. We use $\mathbf{a}\in\mathbb{R}^{d}$ to denote the bias attribute, which is a vector that captures the direction of the bias in the model’s internal representations, and define $\mathbf{r}^{l}$ as the residual at layer $l$.
To estimate the bias attribute, we collect a dataset of standard prompt-image pairs $\mathcal{D}_{\text{standard}}=\{(\mathbf{x}_{i})=(\mathbf{p}_{i},\mathbf{i}_{i})\}_{i=1}^{N_{\text{standard}}}$ and a dataset of prompt-image pairs which elicit biased responses $\mathcal{D}_{\text{bias}}=\{(\mathbf{x}_{i})=(\mathbf{p}_{i},\mathbf{i}_{i})\}_{i=1}^{N_{\text{bias}}}$. Here, $\mathbf{p}_{i}$ represents the text prompt and $\mathbf{i}_{i}$ represents the corresponding image. We compute the activations of the model on both datasets and calculate the difference in means:
@@
Single Step Gradient-based Estimation
By having access to the models weights and a list of words related to a target attribute, we can perform a single step of gradient optimization to create a contrasting set of inputs to the language model. We define $T$ be a list of tokens related to a target attribute.We overload $h^{l}(x,u)$ to represent the output of layer $l$ given the input text $x$ and encoded image tokens $u$.
First, we run a single step of gradient computation on the model to encourage it to mention protected words.
$\mathcal{L}_{\text{bias}}=\frac{1}{|T|}\sum_{t}^{T}||(logit_{t}-(M_{t}))||_{2}^{2}$
where $M_{t}=logit_{t}+max(logits)$. The choice of loss function $L_{bias}$ is discussed and analyzed in section 5.2. Next we compute the gradient w.r.t. the encoded image tokens and simply use the sign of the gradients to modify the image tokens.
@@
where $\epsilon$ is a small constant, and we find computing it based on the average token magnitude to work consistently $\epsilon=\frac{\sigma(|\mathbf{u}|)}{2}$ with $\sigma(\mathbf{u})$ denoting the standard deviation of the image tokens. Then we can simply compute the attribute vector with the modified image representation to compute the linear direction corresponding to the attribute: $\mathbf{a}=\mathbf{h}^{l}(\mathbf{x,u^{\prime}})-\mathbf{h}^{l}(\mathbf{x,u})$
@@
We normalize the bias attribute to have unit length: $\mathbf{a}\leftarrow\mathbf{a}/\|\mathbf{a}\|_{2}$. To ablate the bias attribute, we project the residual at each layer onto the bias attribute and subtract the projection from the residual to get a new residual $\mathbf{r}^{l^{\prime}}=\mathbf{r}^{l}-\mathbf{a}\mathbf{a}^{\top}\mathbf{r}^{l}$. We apply this ablation process to every residual in the LMM, effectively removing the bias attribute direction from the model’s internal representations.
@@ 3.2 Evaluating Steered Models
Following the approach of Howard et al. [17], we evaluate the effectiveness of our steering methods for debiasing through large-scale generation of LMM responses to open-ended prompts. Specifically, given an input image, we generate responses to prompts such as Describe this image in as much detail as possible and What are 5 keywords that describe the characteristics of this person? We use a total of five such prompts inherited from prior work [17] and produce three different LMM responses to each image and prompt by varying the random seed. Our main goal with evaluating our models is to directly measure the effects on bias in an open ended setting.
3.2.1 Datasets
In order to evaluate how open-ended LMM generations vary in response to images depicting people of different social attributes, we utilize images from three datasets which possess social attribute annotations: DA-COCO [10, 44], SocialCounterfactuals [18], and FACET [16]. The SocialCounterfactuals dataset contains 171k synthetic images depicting people of various races, genders, and physical characteristics such as age and body type. Images are grouped into counterfactual sets, in which each image is highly similar in its depiction of a common subject (e.g., a doctor) while varying only according to intersectional social attributes (e.g., race & gender). We use separate subsets of the SocialCounterfactuals dataset for constructing ablation directions and evaluating models for bias related to race, gender, age, and body type. The Demographic Annotations on COCO (DA-COCO) dataset contains 28k images of people with annotations for skin tone and gender, which we use for evaluating bias related to perceived race and gender. Finally, the FACET dataset contains 32,000 diverse, high-resolution images annotated by experts, covering 13 person attributes and 52 person classes. For our evaluation, we utilize the perceived age, race, and gender attribute labels.
3.2.2 Synthetic Image Generation
An essential component of our method is that we can compute steering directions for LMMs from fully synthetic data, that can be successfully applied to real data. For each target attribute e.g. perceived race, age, physical appearance, and gender, we use synthetically generated images from two sources to compute steering directions: the SocialCounterfactuals dataset, and images generated with the Flux-1.0-dev image generator [1]. In the case of SocialCounterfactuals, the images are generated with Stable Diffusion [29], and are constrained to be counterfactual with respect to the target attribute in question e.g. perceived race. For Flux, images align with the target protected attributes but are not constrained to be counterfactuals.

### v2

# Debias your Large Multi-Modal Model at Test-Time via Non-Contrastive Visual Attribute Steering

@@ Abstract
Large Multi-Modal Models (LMMs) have demonstrated impressive capabilities as general-purpose chatbots able to engage in conversations about visual inputs. However, their responses are influenced by societal biases present in their training datasets, leading to undesirable differences in how the model responds when presented with images depicting people of different demographics. In this work, we propose a training-free debiasing framework for LMMs that intervenes on the model’s representations during text generation by constructing a steering vector that reduces reference on protected attributes. Our framework introduces two complementary methods: (1) a dataset-based approach that constructs a steering vector by contrasting model activations on biased and neutral inputs, and (2) a novel optimization-based approach designed for low-resource settings, which constructs the steering vector using a single step of gradient-based perturbation without requiring additional data. Our experiments show that these interventions effectively reduce the propensity of LMMs to generate text related to protected attributes while maintaining sentiment and fluency. Furthermore, we demonstrate that debiased LMMs achieve comparable accuracy to their unmodified counterparts on downstream tasks, indicating that bias mitigation can be achieved without sacrificing model performance.
@@ 1 Introduction
Deep neural networks are well known to exhibit a myriad of societal biases learned from their training datasets [7, 45]. While this effect can be observed to some extent for most models, it is especially prevalent in models trained on web-scale data. Numerous prior works have observed and identified such biases present in modern Large Language Models (LLMs) [4, 8], while other recent work shows that such biases are even more prevalent in Large Multimodal Models (LMMs)  [^1] LMMs refer specifically to generative multimodal models. We refer to the general class of vision-language models as VLMs [6] such as LLaVA-1.5 [24], or Llama-3.2 [11] that integrate a vision backbone with a pretrained LLM. Given that LLMs are often pretrained on relatively uncurated web-scale data [30], any further integration with additional models will inherit the particular biases of the LLM. Without additional safety tuning, these pre-existing biases may be amplified further when an LLM is augmented with pretrained visual capabilities [17], which also come with a distinct set of implicit societal biases present in the visual pretraining data. Evaluating and mitigating potentially harmful behaviors induced by these societal biases is becoming increasingly important in order to safely deploy multimodal generative AI systems that utilize LMMs.
In this work, we address a form of such biases by seeking to reduce the propensity of an LMM to generate protected attribute-related text unnecessarily. While it is often benign and even useful to make use of protected attributes in generation, LMMs have been found to suffer from group-dependent sentiment disparities [27] as well as conditional inaccuracies given certain protected attributes [18, 44]. In many real-world applications, such group-conditional performance differences are undesirable. Therefore, our goal is to provide methods for mitigating this behavior via modification of the internal representations of an LMM with respect to a set of target protected attributes.
Figure 1: Our proposed approaches for debiasing LMMs. Left: our PARzero approach solves for a steering vector with a single optimization step toward the target attributes, requiring no additional data. Middle: our PAR approach constructs a steering vector from a dataset of contrasting activations given by attribute-related Q&A. Right: LMMs with basic debiasing mechanisms like reprompting are not effective for debiasing on real world data.
@@
A variety of methods have been proposed recently for debiasing LLMs and discriminative VLMs individually [22, 32]. In contrast, relatively little prior work has focused specifically on debiasing LMMs. One naive approach could be to simply define a list of words or symbols that are prohibited from being generated at inference time. However, a more nuanced approach is required for attributes with many or ambiguous meanings, such as perceived race (e.g., black, white) [^2] Wherever we use words such as race or gender throughout this work, we mean perceived race and perceived gender. The labels we used for such social attributes are inherited from prior work and are not a reflection of our own value judgments. See Section 6 for additional discussion.. Many other existing debiasing approaches for LLMs and VLMs focus on fine-tuning models with additional curated data to reduce bias. However, training is a blunt tool that often results in undesirable outcomes such as a degradation in task-specific performance. Fine-tuning is also labor and computationally intensive, requiring the collection of an additional (likely large-scale) dataset that can appropriately debias the model. Despite prior efforts [17], there remains no canonical recipe for constructing a debiasing dataset large enough to meet data requirements for training LMMs while also containing fine-grained social attribute annotations. Our work addresses these limitations by introducing a training-free approach for debiasing LMMs which can be applied to any attribute at inference time without degrading model performance.
Specifically, we propose a novel method for identifying and ablating (removing) biased representations within LMMs. Recent work has shown that LLMs tend to encode high-level concepts as linear directions in feature space [35], making it possible to steer a specific feature’s contribution to the model’s output. For example, adding a “steering vector” can amplify a desired feature [47], whereas ablating that direction can remove the feature [3]. We introduce two methods for constructing steering vectors: (1) a dataset-based approach, where steering vectors are computed offline by comparing model activations from contrasting datasets—one that provokes the feature of interest and one that does not, and (2) a novel optimization-based approach, which constructs the steering vector at inference time using just the input image and the target attribute for removal.
@@
Our methods differ from prior LLM steering work in two fundamental ways. First, our dataset-based method conditions the LMM on biased multimodal inputs to induce the target representations. Second, when no data is available, our optimization-based approach computes a steering vector at inference time using just the input image and the target attribute for removal. Instead of crafting a dataset of contrasting prompts, we only need to leverage the information already available to the LMM, the input image. Specifically, during inference we perform a single step of gradient descent on the visual representation in the direction of the target attribute(s). For both approaches, steering is performed by extracting downstream activations from a single layer that follow from both the perturbed and the unchanged visual representations. The normalized difference between them is used as the steering direction.
@@
  • We propose an two training-free algorithms to find effective steering directions for LMMs. PAR (Protected Attribute Removal), that uses a dataset of biased inputs, and a novel method PARzero designed for low-resource settings. PARzero uses no extra data, and steers by performing a single optimization step on the input visual representation.
  • We show that our proposed approaches can effectively debias two popular LMM model families LLaVA and Llama3. We reduce their propensity to generate text related to protected attributes on the SocialCounterfactuals, COCO, and FACET datasets. We find that our results hold across perceived race, physical appearance, age, and gender attributes.
  • We analyze the steered and unsteered generations across all datasets and find that steering produces equitable group-level sentiment without degrading accuracy or language modeling capabilities.
@@ 2 Related Work
Model steering and interventions. Prior work on controlling LLM outputs has explored interventions on model inputs, outputs, or intermediate representations. Modifying prompts to induce or ablate a feature [41, 40] is one such approach, but is commonly an iterative and inefficient method. Intervening on outputs has been explored via token-level block lists [14], domain classifiers [21] to trigger regeneration, or decoding strategies that decrease probabilities of generating certain tokens [29]. Intervening on intermediate model representations is the most related approach to our work. This often entails the discovery of interpretable features that not only explain model behavior [47, 19], but can also be used to intervene and steer the model towards outputs with certain characteristics or content.
Templeton et al. [35] showed that interpretable features can be found by training a sparse autoencoder [19] to decompose the activations of an LLM into separable features. These features were found to be monosemantic – able to locally augment downstream outputs when manually introduced during inference. Turner et al. [37] showed that a steering vector for amplifying behavior in LLMs could be constructed with a single pair of contrasting prompts, and Arditi et al. [3] demonstrated that ablation of an abstract concept like refusal in LLMs can be performed via dataset-based steering.
@@
While several approaches have been proposed for mitigating social biases in VLMs [39, 5, 43, 31, 10, 33, 18], prior research on addressing such biases in LMMs is lacking. Sathe et al. [28] and Fraser and Kiritchenko [12] utilized synthetically generated images to analyze the presence of bias in LMMs, but do not address bias mitigation strategies. Hinck et al. [16] employ steering-based strategies for improving the linguistic fidelity of LMMs, but do not consider other kinds of bias. Howard et al. [17] also leveraged synthetic images from the SocialCounterfactuals dataset [18] to measure bias in LMMs but at a much larger scale, finding that LMMs possess more bias than the corresponding LLM from which they were trained. While representation-based steering for reducing societal biases has been demonstrated in LLMs such as Claude 3 [35], our work is the first to demonstrate successful inference-time representation steering for reducing bias in LMMs.
Figure 2: PARzero’s gradient-based steering direction computation. (Step 1) Given an input image, prompt the LMM to produce text related to the target attribute and store the intermediate activations. (Step 2) Compute modified image tokens $\mathbf{u^{\prime}}$ by performing a single optimization on the image tokens to maximize the downstream logits $\mathbf{z}$ corresponding to the target token indices, where $\bar{\mathbf{z}}$ is a copy of $\mathbf{z}$ with amplified target tokens. (Step 3) Recompute the activations of the LLM with the modified image tokens $\mathbf{u^{\prime}}$ and compute attribute direction $a$ as the difference between the two stored activation vectors. (Step 4) Complete the query while removing $a$ from all intermediate layers of the LLM.
@@ 3 Methodology
Motivated by prior work on steering LLM behavior via intervening on representations, we propose an approach for debiasing LMMs by identifying and ablating a linear direction in the LMM’s activation space that corresponds to its propensity to mention a set of target protected attributes. We propose two methods for achieving protected attribute removal: PAR and PARzero. PAR leverages a synthetic dataset of biased and neutral inputs to induce the target representations. For low resource environments, or for increased flexibility, PARzero requires no additional data, and steers by taking a single step of gradient descent toward the target attribute. Both approaches compute a steering vector by taking the difference between internal representations when the model is generating text with and without protected attribute mentions.
@@ 3.1 Steering Direction Estimation
PAR
Let $\mathcal{M}$ denote an LMM, and $\mathbf{h}^{l}\in\mathbb{R}^{d}$ represent the activations at a layer of interest $l$, where $d$ is the dimensionality of the hidden state. We use $\mathbf{a}\in\mathbb{R}^{d}$ to denote the steering vector to be computed; $\mathbf{a}$ captures the direction of the relevant bias in the internal representation at layer $l$.
To estimate the bias attribute, we collect a dataset of inputs (image-prompt pairs) that elicit biased responses with repsect to a target attribute $\mathcal{D}_{\text{bias}}=\{(\mathbf{x}_{i})=(\mathbf{p}_{i},\mathbf{i}_{i})\}_{i=1}^{N_{\text{bias}}}$, as well as a dataset of arbitrary inputs $\mathcal{D}_{\text{standard}}=\{(\mathbf{x}_{i})=(\mathbf{p}_{i},\mathbf{i}_{i})\}_{i=1}^{N_{\text{standard}}}$, where $\mathbf{p}_{i}$ is the prompt and $\mathbf{i}_{i}$ refers to the corresponding synthetic image. See section (13) for details of our synthetic image generation pipeline. We compute the activations of the model on both datasets and calculate the difference in means:
@@
PARzero
Given a list of target attribute-related token indices $T$, and access to the intermediate visual representation of the LMM [^3] For LLaVA we perturb the visual tokens at the input to the LLM. For Llama we perturb the visual features at the multimodal projector., our goal is to modify the visual representation such that we induce sufficient contrast in the downstream activations of layer $l$ to construct a steering vector (see Figure (2)). Let $h^{l}(x,\mathbf{u})$ be the output of LLM layer $l$ given the prompt $x$ and encoded image tokens $\mathbf{u}$. We wish to amplify the output logits $\mathbf{z}$ for target tokens indices $t\in T$, which we accomplish via a single-step perturbation of $\mathbf{u}$. To find a perturbation direction we construct the following loss:
$\mathcal{L}_{\text{bias}}=||(\mathbf{z}-\bar{\mathbf{z}})||_{2}^{2}$
where $\bar{\mathbf{z}}$ is a rescaling of the target token indices:
$\bar{z}_{i}=\begin{cases}z_{i},&\text{if }i\in T,\\[6.0pt] z_{i}+\max(\mathbf{z}),&\text{o.w.},\end{cases}\quad\text{for }i=1,\ldots,\lvert\mathbf{z}\rvert.$
We analyze this loss function $L_{bias}$ in section (7) of the supplementary material. We make use of the sign of the gradient $\nabla_{\mathbf{u}}\mathcal{L}_{\text{bias}}$ to modify the image tokens as follows:
@@
where $\epsilon=\frac{\sigma(|\mathbf{u}|)}{2}$ with $\sigma(\mathbf{u})$ is a small constant based on the average token magnitude [^4] We set $\epsilon$ dynamically given that attribute-related tokens probabilities may differ widely at inference time.. We then simply compute the steering vector with the modified image representation, yielding the linear direction corresponding to the target attribute: $\mathbf{a}=\mathbf{h}^{l}(\mathbf{x,u^{\prime}})-\mathbf{h}^{l}(\mathbf{x,u})$
@@
For both PAR and PARzero, we normalize the steering vector to have unit length: $\mathbf{a}\leftarrow\frac{\mathbf{a}}{\|\mathbf{a}\|_{2}}.$ Steering is performed for each forward pass; we project the hidden state activations at each layer onto the steering vector and remove the projected component to obtain the debiased representation i.e. $\mathbf{h}^{l^{\prime}}=\mathbf{h}^{l}-\text{proj}_{\mathbf{a}}(\mathbf{h}^{l})$, where $\text{proj}_{\mathbf{a}}(\mathbf{h}^{l})=(\mathbf{a}^{\top}\mathbf{h}^{l})\mathbf{a}.$ We apply this feature ablation process to all layers of the LMM language model, effectively removing the contribution of the bias attribute direction from the model’s internal representations.
@@ 3.2 Evaluating Steered Models
Following the approach of Howard et al. [17], we evaluate the effectiveness of our steering methods for debiasing through large-scale generation of LMM responses to open-ended prompts. Specifically, given an input image, we generate responses to prompts such as Describe this image in as much detail as possible and What are 5 keywords that describe the characteristics of this person? We use a total of five such prompts inherited from prior work [17] and produce three different LMM responses to each image and prompt by varying the random seed. Our main goal with evaluating our models is to directly measure the effects on bias in an open ended setting. See section (12) of the supplementary material for analysis on less-open ended evaluations and section (8) complete details of our experimental settings.
3.2.1 Evaluation datasets
To evaluate how open-ended LMM generations vary in response to images depicting people of different social attributes, we utilize images from three datasets which possess social attribute annotations: SocialCounterfactuals [18], DA-COCO [9, 44], and FACET [15]. We pair each dataset with our open-ended prompts to evaluate our approach. The SocialCounterfactuals dataset contains 171k synthetic images depicting people of various races, genders, and physical characteristics such as age and body type. Images are grouped into counterfactual sets, in which each image is highly similar in its depiction of a common subject (e.g., a doctor) while varying only according to intersectional social attributes (e.g., race & gender). We use subsets of the SocialCounterfactuals dataset for constructing steering directions and evaluating models for bias related to race, gender, age, and body type. The Demographic Annotations on COCO (DA-COCO) dataset contains 28k images of people with annotations for skin tone and gender, which we use for evaluating bias related to perceived race and gender. Finally, the FACET dataset contains 32k diverse, high-resolution images annotated by experts, covering 13 person attributes and 52 person classes. For our evaluation, we utilize the perceived age, race, and gender attribute labels. See section (9) of the supplementary material for additional evaluation details.